# Out of Length Text Recognition with Sub-String Matching

### v2

@@ 3.3 Comparison with State-of-the-arts
Results on English Benchmarks. We apply SMTR to Common benchmarks, Union14M benchmarks and LTB, and compare it with existing STR models. The results are in Tab. 4. Compared with PARSeq, one leading method in short text benchmarks, SMTR exhibits quite similar accuracy. In addition, it correctly recognizes half of LTB text, which is incapable for PARSeq. Note that AR-STR also achieves very competitive results on short text benchmarks. However, its weak extrapolation capability restricts the performance on LTB. On the other hand, FocalSVTR performs decently on short text benchmarks, but on LTB it performs on par with LISTER, which is a dedicated model for long text recognition. The results above are in line with the observation in Fig. 1. To sum, by introducing a novel sub-string matching-based recognition paradigm, SMTR well recognizes both short and long text, which would be a promising property for many STR-related applications.
@@
Results on Chinese Benchmarks. The OOL challenge is pervasive across languages. To verify the multilingual adaptability of SMTR, we conduct evaluations on the Chinese text recognition benchmark (CTR) (Chen et al. 2021). In line with CCR-CLIP (Yu et al. 2023), SMTR refrains from utilizing data augmentation (w/o Aug) to ensure a fair comparison. As presented in Tab. 5, SMTR outperforms LISTER by 6.83% on $Scene$ subset. Note that SMTR gets this result without training on long text while LISTER has no imposed length limitation during training. Despite this discrepancy, SMTR still maintains an accuracy gain of 10.61% over LISTER in Chinese long text, as depicted in Tab. 6. Meanwhile, SMTR achieves a new state-of-the-art on CTR, boosting an accuracy gain of 3.53% compared to CAM (Yang et al. 2024), the previous best method. These results demonstrate SMTR’s great adaptability in multilingual recognition.
Method $Scene$ $Web$ $Doc$ $HW$ Avg
@@
CRNN (2017) 53.4 57.0 96.6 50.8 64.45 12.4
ASTER (2019) 61.3 51.7 96.2 37.0 61.55 27.2
MORAN (2019) 54.6 31.5 86.1 16.2 47.10 28.5
SAR (2019) 59.7 58.0 95.7 36.5 62.48 27.8
SEED (2020) 44.7 28.1 91.4 21.0 46.30 36.1
MASTER (2021) 62.8 52.1 84.4 26.9 56.55 62.8
ABINet (2021) 66.6 63.2 98.2 53.1 70.28 53.1
TransOCR (2021) 71.3 64.8 97.1 53.0 71.55 83.9
SVTR-B (2022) 71.7 73.8 98.2 52.2 73.98 26.3
CCR-CLIP (2023) 71.3 69.2 98.3 60.3 74.78 62.0
LISTER (2023) 73.0 - - - - 55.0
DCTC (2024) 73.9 68.5 99.4 51.0 73.20 40.8
CAM (2024) 76.0 69.3 98.1 59.2 76.80 135
SMTR w/o Aug 79.8 80.6 99.1 61.9 80.33 20.8
SMTR w/ Aug 83.4 83.0 99.3 65.1 82.68 20.8
Table 5: Results of different models on CTR dataset.
Method $L_{\leq 25}$ $L_{\textgreater 25}$ Weighted Avg Arithmetic Avg
63136 510
LISTER (2023) 73.00 35.00 72.69 54.00
SMTR 79.83 45.61 79.56 62.74
Table 6: Short and long text recognition results on $Scene$.
@@ 4 Conclusion
Appendix A Related Work
Existing STR methods mainly focus on word-level text recognition. Among them, attention-based (Lee and Osindero 2016) methods are intensively studied for their impressive performance. Some of these methods (Bautista and Atienza 2022; Yue et al. 2020; Wang et al. 2021; Du et al. 2023) use learnable position embeddings to accurately learn context information. However, because learnable position embeddings are typically not scalable, these methods can only recognize short text. In contract, methods (Shi et al. 2019; Li et al. 2019; Sheng, Chen, and Xu 2019; Fang et al. 2021; Zhang et al. 2023; Zheng et al. 2024; Du et al. 2024) employ LSTM (Hochreiter and Schmidhuber 1997; Chung et al. 2014) or masked self-attention mechanism with sinusoidal positional encoding (Lee and Osindero 2016) to model context. They have some length extrapolation capability. Nevertheless, these models are primarily trained on short text. Therefore their inference ability is limited when dealing with long text, resulting in a significant decrease in recognition accuracy. Additionally, some methods do not rely on well-designed attention-based decoders, such as CTC-based (Graves et al. 2006) methods (Shi, Bai, and Yao 2017; Hu et al. 2020; Du et al. 2022) and length-insensitive LISTER (Cheng et al. 2023). They do not perform well on short text recognition although exhibiting slightly stronger capability in long text recognition.
Appendix B Pseudo-code of Inference Augmentation
Algorithm 2 presents the implementation of inference augmentation in Python style. Specifically, the Slice operation splits an image into three sub-images as shown in Fig. 4. It is then followed by bi-directional inference using Inference to obtain the recognition results of Img1 and Img2. In the following, the third Inference is performed on Img3 with sub-strings from both sides given, which can correct mis-recognition caused by Slice. Finally, the full long text recognition result is obtained with a straightforward post-processing. The operation Inference is detailed in Algorithm 1.
Appendix C Details of the Sub-String Matcher Module
@@
Method Common Benchmarks Union14M-L Benchmarks LTB Params (×$\times 10^{6}$)
@@
Table 9: Results of SMTR and existing models on short text benchmarks and LTB when trained on synthetic datasets. * represents that the result is evaluated on Union14M-L benchmarks using the model they released.
Appendix D Results when trained on synthetic datasets
@@
Appendix E Visualizations of attention maps
@@
Appendix F Bad cases of SMTR

### v3

@@ 4.3 Comparison with State-of-the-arts
Results on English Benchmarks. We apply SMTR to Common benchmarks, Union14M-Benchmark and LTB, and compare it with existing STR models. The results are in Tab. 4. Compared with PARSeq, one leading method in short text benchmarks, SMTR exhibits quite similar accuracy. In addition, it correctly recognizes half of LTB text, which is incapable for PARSeq. Note that AR-STR also achieves very competitive results on short text benchmarks. However, its weak extrapolation capability restricts the performance on LTB. On the other hand, FocalSVTR performs decently on short text benchmarks, but on LTB it performs on par with LISTER, which is a dedicated model for long text recognition. The results above are in line with the observation in Fig. 1. To sum, by introducing a novel sub-string matching-based recognition paradigm, SMTR well recognizes both short and long text, which would be a promising property for many STR-related applications.
@@
Results on Chinese Benchmarks. The OOL challenge is pervasive across languages. To verify the multilingual adaptability of SMTR, we conduct evaluations on the Chinese text recognition benchmark (CTR) (Chen et al. 2021). In line with CCR-CLIP (Yu et al. 2023), SMTR refrains from utilizing data augmentation (w/o Aug) to ensure a fair comparison. As presented in Tab. 5, SMTR outperforms LISTER by 6.83% on Scene subset. Note that SMTR gets this result without training on long text while LISTER has no imposed length limitation during training. Despite this discrepancy, SMTR still maintains an accuracy gain of 10.61% over LISTER in Chinese long text ($L_{>25}$). Meanwhile, SMTR achieves a new state-of-the-art on CTR, boosting an accuracy gain of 3.53% compared to CAM (Yang et al. 2024), the previous best method. These results demonstrate SMTR’s great adaptability in multilingual recognition.
Method Scene Web Doc HW Avg $L_{>25}$
@@
CRNN (2017) 53.4 57.0 96.6 50.8 64.45 - 12.4
ASTER (2019) 61.3 51.7 96.2 37.0 61.55 - 27.2
MORAN (2019) 54.6 31.5 86.1 16.2 47.10 - 28.5
SAR (2019) 59.7 58.0 95.7 36.5 62.48 - 27.8
SEED (2020) 44.7 28.1 91.4 21.0 46.30 - 36.1
MASTER (2021) 62.8 52.1 84.4 26.9 56.55 - 62.8
ABINet (2021) 66.6 63.2 98.2 53.1 70.28 0.0 53.1
TransOCR (2021) 71.3 64.8 97.1 53.0 71.55 - 83.9
SVTR-B (2022) 71.7 73.8 98.2 52.2 73.98 - 26.3
CCR-CLIP (2023) 71.3 69.2 98.3 60.3 74.78 - 62.0
LISTER (2023) 73.0 - - - - 35.00 55.0
DCTC (2024b) 73.9 68.5 99.4 51.0 73.20 - 40.8
CAM (2024) 76.0 69.3 98.1 59.2 76.80 - 135
SMTR w/o Aug 79.8 80.6 99.1 61.9 80.33 45.61 20.8
SMTR w/ Aug 83.4 83.0 99.3 65.1 82.68 49.41 20.8
Table 5: Results of different models on CTR dataset. $L_{>25}$ means long text recognition results on Scene.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China (No. 2022YFB3104703) and in part by the National Natural Science Foundation of China (No. 32341012).
Appendix A Pseudo-code of Inference Augmentation
Algorithm 2 presents the implementation of inference augmentation (IA) in Python style. Specifically, the Slice operation splits an image into three sub-images as shown in Fig. 4. It is then followed by bi-directional inference using Inference to obtain the recognition results of Img1 and Img2. In the following, the third Inference is performed on Img3 with sub-strings from both sides given, which can correct mis-recognition caused by Slice. Finally, the full long text recognition result is obtained with a straightforward post-processing. The operation Inference is detailed in Algorithm 1.
By leveraging IA, SMTR not only achieves enhanced recognition accuracy for long texts but also significantly accelerates inference speed. As shown in Fig. 8, IA improves SMTR’s efficiency in long text recognition. This improvement is primarily attributed to IA’s split-merge processing mechanism, which substantially reduces the number of iterations required for recognizing lengthy texts.
Appendix B Details of the Sub-String Matcher Module
@@
Method Common Benchmarks Union14M-Benchmark LTB Params (×$\times 10^{6}$)
@@
Table 8: Results of SMTR and existing models on short text benchmarks and LTB when trained on synthetic datasets. * represents that the result is evaluated on Union14M-Benchmark using the model they released.
Appendix C Results when trained on synthetic datasets
@@
Appendix D Visualizations of attention maps
@@
Appendix E Bad cases of SMTR